Find the volume of the ellipsoid x^2+y^2+5z^2=64.
1
Expert's answer
2020-11-17T16:14:18-0500

"x^2 +y^2+5z^2 = 64"

"\dfrac{x^2}{8^2} +\dfrac{y^2}{8^2} +\dfrac{z^2}{(\cfrac{8}{\sqrt5})^2} = 1\\
a = 8;\space b =8;\space c = {\cfrac{8}{\sqrt5}}\\
V =\cfrac{4}{3}\pi*a*b*c"

"V = \cfrac{4}{3}\pi*8*8*\dfrac{8}{\sqrt5} = \cfrac{512\sqrt5}{15}\pi =76.325\pi"